Snookers Lumber can convert logs into lumber or plywood. In a given day, the mill turns twice as many units of plywwod as lumber. It makes a profit of $20 on a unit of lumber and $50 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $12,480?
Snookers Lumber must produce and sell ___ units of lumber and ___ units of plywood to make a profit of $12,480.

Thank you for your help :)

Answer by rfer(16322)

50(2L)+20L=12480
100L+20L=12480
120L=12480
L=104
P=208
104*20+208*50=12480